# Document Type Classification using File Names

### v1

@@ 6 Experiments
SVM, Naives Bayes, Random Forest, XGBoost, and BERT classifiers are trained on only the indicative file names from the Web Search dataset, and evaluated against the rest of the data from the Web Search dataset and the Common Crawl dataset. In an additional experimental set up, all the classifiers are trained on indicative file names and ambiguous file names (the latter being a single additional classification class) from the Web Search dataset and evaluated against both datasets. We perform 5-fold cross validation and calculate the average prediction accuracy and prediction rate of different setups (defined in Section 4.3). The subsections below discuss the accuracy and speed of our file name classifiers, as well as their ability to distinguish ambiguous data.
@@ 6.1 Model Accuracy for Indicative File Names
First, we show that our approach can accurately classify indicative file names. Table 2 shows the confidence thresholds and tokenizer combination that maximizes the prediction accuracy of every model while the prediction rate is above 0.8 when tested on only indicative data. Predictions that result in a confidence score lower than the threshold or produce an ‘ambiguous’ label are deferred by the file name classifier. With a prediction rate of 0.805, the Random Forest file name classifier with a Trie tokenizer trained with a negative class produces the highest prediction accuracy of 99.6% showing that this file name classifier can produce accurate predictions for the majority of the indicative data while deferring the rest. SVM, Random Forest, XGBoost, and BERT when trained with a negative class, all achieved above 99% accuracy with a prediction rate above 0.8. When the negative class is not used in training, every model’s prediction accuracy decreases substantially, with the two most accurate classifiers with BERT and Random Forest, achieved 94.8% and 94.1% accuracy, respectively. We observed insubstantial differences in prediction accuracy for every file name classifier with different tokenizers.
Figure 3: The confidence score distribution for Random Forest classifier with Trie tokenizer on indicative and ambiguous file names from the Web Search dataset.
@@
Models Without Negative Class With Negative Class
Tokenizer Threshold Accuracy Tokenizer Threshold Accuracy
SVM Trie 0.60 0.748 BERT 0.46 0.931
Naive Bayes BERT 0.59 0.729 BERT 0.29 0.782
Random Forest BERT 0.61 0.934 BERT 0.65 0.967
XGBoost Trie 0.83 0.879 BERT 0.71 0.915
BERT Trie 0.97 0.866 BERT 0.97 0.842
Table 4: The confidence threshold value and tokenizer that maximizes the prediction accuracy for every model trained with or without a negative class while the prediction rate is above 0.178 for all confidence threshold between 0 and 1 incremented by 0.01. The models are trained on the Web Search dataset and evaluated on the Common Crawl dataset.
6.2 Detection Ability for Ambiguous File Names
We evaluate our file name classifiers against the ambiguous file names and evaluate our model’s robustness predicting ambiguous file names by calculating the AUROC of the confidence scores between the indicative and ambiguous file names. Figure 3 shows minor overlap between the confidence scores for the indicative and ambiguous file names for the Random Forest file name classifier with the Trie tokenizer.
Table 3 displays the AUROC of the confidence scores between indicative and ambiguous file names for all models and each tokenizer. The Random Forest file name classifier with the Trie tokenizer achieves the highest AUROC which shows that this classifier produces confidence scores that best differentiate indicative and ambiguous file names. The differences in AUROC for different tokenizers are also not substantial for all models.
Due to the differences in how we obtain confidence scores for each models, the optimal confidence threshold differs between models. Increasing the confidence threshold increases the prediction accuracy while decreasing the prediction rate. In practice, we must select an appropriate confidence threshold to balance the tradeoff between the prediction accuracy and the fraction of documents accurately predicted as indicative — the documents that are deferred as ambiguous would require further examination using a deeper and more complex model. Thus, the average time and resource expenditure per document classified would increase with the increase in confidence threshold for predicting a document file name as indicative.
Figure 4 shows the prediction accuracy and prediction rate of the Random Forest classifier on all the file names in the Web Search dataset (both indicative and ambiguous) for for all confidence thresholds between 0 and 1, in intervals of 0.01 per iteration. Figure 5 shows the overall accuracy and overall time per document required for a Random Forest classifier with the Trie tokenizer trained with a negative class and a DiT model to process all the deferred predictions from the file name classifier tested on the Web Search data including indicative and ambiguous file names. With the assumption that all DiT predictions are correct, overall accuracy is calculated by prediction accuracy $\times$ prediction rate $+$ ($1-$ prediction rate), and overall time is calculated by average inference time of Random Forest + average inference time of DiT * ($1-$ prediction rate). The inference time for each model is discussed in Section 6.4. At a confidence threshold of 0.66, the improvement in prediction accuracy begins to taper off, while the decrease in prediction rate accelerates, and the overall time increases more steeply. This indicates that 0.66 is an optimal confidence threshold for the Random Forest file name classifier.
Figure 4: Prediction Accuracy vs Prediction Rate at all confidence score threshold between 0 and 1 in increments of 0.01 for a Random Forest classifier with the Trie tokenizer trained with a negative class predicted on all the file names in the Web Search dataset including ambiguous file names.
Figure 5: Overall Accuracy vs Overall Time (seconds) at all confidence score threshold between 0 and 1, incremented by 0.01 for a Random Forest classifier with the Trie tokenizer trained with a negative class predicted on all the file names in the Web Search dataset including ambiguous file names, and DiT model to process the documents deferred by the Random Forest model.
6.3 Evaluation on Common Crawl Dataset
We train file name classifiers on the Web Search dataset described in Section 5.1 and evaluate our models against the Common Crawl dataset described in Section 5.2.
Recall that 22.2% of the file names in our Common Crawl dataset belong to a category in the Web Search dataset. Without further annotating whether each of those in-scope file names are indicative, we evaluate how accurate our file name classifiers are while predicting at least 80% of the in-scope data, which is 17.8% of all the documents. Table 4 displays the confidence threshold value and the tokenizer that maximizes the prediction accuracy while achieving a prediction rate of at least 0.178. The result displays that the Random Forest classifier with a BERT tokenizer achieves 96.7% prediction accuracy when trained with a negative class, and is the most robust classifier at predicting the Common Crawl dataset. The SVM file name classifier trained with a negative class and the Random Forest classifier trained without a negative class also reached above 93% prediction accuracy with the BERT tokenizer.
The results indicate that the Random Forest file name classifier trained with a negative class can accurately predict the majority of the in-scope documents from a distinctive dataset with a large portion of out-of-scope documents.
@@ 6.3 Evaluation on Common Crawl Dataset
6.4 Inference Time
Among all the models, Naive Bayes was the fastest, with a average log inference time of $-13.34$, requiring $1.61\times 10^{-6}$ seconds per prediction. The BERT file name classifier’s inference time is substantially slower than all four other models, requiring $6.14\times 10^{-2}$ seconds per prediction. The Random Forest classifier that achieved the highest prediction accuracy on both the Web Search dataset and the Common Crawl dataset requires $1.30\times 10^{-4}$ seconds per prediction.
In comparison, when used on the same computer as the file name classifiers, the DiT model on average requires $0.0574$ seconds per image, and the BERT model that uses the first 512 tokens of a document as input requires $0.0678$ seconds per document, both by analyzing the content of the documents instead of the file names. The robust Random Forest file name classifier is 442.43x faster than the DiT model even if the DiT model only processes the first page of every document and is 552.60x faster than the BERT model that processes the content of the documents, offering a dramatic speedup.
@@ 6.5 Trade-offs for $k$ Value Selection
Figure 6: Prediction Accuracy vs Prediction Rate at 0.8 confidence threshold for the Random Forest file name classifier with a BERT tokenizer using various k-values on indicative file names from the Web Search dataset.
In Section 4.1, we mentioned that tokens with TF-IDF larger than $k$ are considered "keywords" and added to the tokenizer. We examine the trade-offs for the selection of $k$ value for the BERT tokenizer by comparing the prediction accuracy and prediction rate with various $k$ values. The previous subsections indicate that the Random Forest file name classifier produces the highest prediction accuracy with a fixed minimum prediction rate. Figure 6 displays the prediction accuracy and prediction rate at 0.8 confidence threshold for the Random Forest file name classifiers with $k$ values from 0 to 0.5. For $k$ values between 0.04 and 0.43, the prediction rate remains steady around 0.865 while the prediction accuracy is the highest for $k$ values between 0.025 to 0.132. Thus, a $k$ value should be selected between 0.025 and 0.132.
@@ 7 Conclusion
We present a new approach to multi-class document classification that dramatically improves classification efficiency while achieving high accuracy. Our evaluation shows that a Random Forest classifier, when tasked to predict at least 80% of the in-scope data, can achieve 99.6% prediction accuracy on indicative file names when trained and tested on datasets with the same category distribution with a Trie tokenizer. It also achieve 96.7 % prediction on the in-scope file names when trained and tested on data with different distribution of categories with a BERT tokenizer. The Random Forest classifier with a Trie tokenizer can also produce confidence scores that can effectively distinguish indicative and ambiguous file names, resulting in an AUROC of 0.93 for confidence scores between indicative and ambiguous. Meanwhile, the Random Forest classifier is 442.43x faster than the DiT model, indicating that uses the Random Forest file name classifier can provide a substantial overall speedup in document classification. We hope that our technique and analysis will enable faster and computationally efficient document classifiers without sacrificing on accuracy.
Future Work
In this study, we solely focused on document classification using file names and obtained promising results. We believe file metadata — such as file size, page count, number of images, etc. — can be used along with file names to improve the classifier accuracy without substantial increase in inference time. Future work can investigate the potential benefits of combining metadata with file name to improve classification accuracy and or handle ambiguous data or out-of-scope data more effectively.
@@ Appendix A Appendix
We provide both the Web Search and Common Crawl datasets in an anonymous Github repository: https://github.com/frank7li/Document-Type-Classification-using-File-Names.
@@
Table 6: Document type categories used in the Web Search dataset.

### v2

@@ 6 Results
SVM, Naives Bayes, Random Forest, XGBoost, and BERT classifiers are trained on 80% of the indicative file names from the Web Search dataset, and evaluated against the rest of the data from the Web Search dataset and the entire Common Crawl dataset. We perform 5-fold cross validation and calculate the average accuracy, prediction accuracy, and prediction rate (defined in Section 4.3). The subsections below discuss the accuracy and speed of our file name classifiers, as well as their ability to distinguish ambiguous and out-of-scope data.
@@ 6.1 Model Accuracy for Indicative File Names
First, we show that our approach can accurately classify indicative file names. Table 2 and table 3 shows the accuracy for all models with the trie and BERT tokenizer on the Web Search dataset respectively. The BERT file name classifier achieves the highest accuracy of 95.51 with a trie tokenizer, while Random Forest, Naive Bayes, and SVM also achieves above 94% accuracy.
Additionally, we show the prediction accuracy with at least 0.9 prediction rate iterating through all confidence thresholds between 1 and 0, decrementing by 0.01. Predictions that result in a confidence score lower than the threshold are deferred by the file name classifier. Deferring less than 10% of the data, the Random Forest classifier with the BERT tokenizer produces the highest accuracy of 99.78% among all combinations of models and tokenizers. SVM and BERT also achieved more than 99% accuracy while predicting more than 90% of the data, while there is a gap in accuracy for the Naive Bayes model.
In addition, we display the performance of the five models trained on the Web Search dataset and tested on the in-scope, indicative file names from the Common Crawl dataset in table 4 and table 5. The Random Forest classifier with the trie tokenizer achieves 96.57% prediction accuracy with above 0.9 prediction rate. We observe that the Random Forest classifier obtains the highest prediction accuracy among all five models, consistent with the evaluation results from the Web Search dataset.
We observe insubstantial differences in accuracy and prediction accuracy for every file name classifier with different tokenizers.
Models Accuracy Threshold Prediction Accuracy
SVM 94.06 0.55 99.34
Naive Bayes 94.13 0.37 98.02
Random Forest 94.92 0.51 99.63
XGBoost 92.61 0.73 99.41
BERT 95.51 0.96 99.42
Table 2: The accuracy and prediction accuracy with prediction rate more than 0.9 for all models with the trie tokenizer on the Web Search dataset.
Models Accuracy Threshold Prediction Accuracy
SVM 94.06 0.49 99.12
Naive Bayes 94.39 0.35 98.17
Random Forest 94.72 0.50 99.78
XGBoost 90.76 0.6 98.97
BERT 94.92 0.97 99.35
Table 3: The accuracy and prediction accuracy with prediction rate more than 0.9 for all models with the BERT tokenizer on the Web Search dataset.
Models Accuracy Threshold Prediction Accuracy
SVM 84.23 0.22 92.24
Naive Bayes 87.96 0.25 95.00
Random Forest 88.24 0.31 96.57
XGBoost 84.93 0.30 93.69
BERT 83.45 0.77 89.70
Table 4: The accuracy and prediction accuracy with prediction rate more than 0.9 for all models with the trie tokenizer on the indicative documents in the Common Crawl dataset.
Models Accuracy Threshold Prediction Accuracy
SVM 84.37 0.23 92.82
Naive Bayes 86.48 0.26 92.30
Random Forest 87.82 0.26 96.33
XGBoost 84.86 0.32 93.75
BERT 82.96 0.70 89.31
Table 5: The accuracy and prediction accuracy with prediction rate more than 0.9 for all models with the BERT tokenizer on the indicative documents in the Common Crawl dataset.
6.2 Detection Ability for Ambiguous and Out-of-Scope File Names
We train file name classifiers on the Web Search dataset described in Section 5.1 and evaluate our models against the Common Crawl dataset described in Section 5.2. We test our file name classifiers against the ambiguous and out-of-scope file names and evaluate our model’s robustness predicting ambiguous file names by calculating the AUROC of the confidence scores between the indicative and ambiguous file names.
Figure 3 shows minor overlap between the confidence scores for the indicative and ambiguous file names for the Random Forest file name classifier with the trie tokenizer.
Table 6 displays the AUROC of the confidence scores between indicative and ambiguous file names for all models and each tokenizer. The Random Forest file name classifier with the trie tokenizer achieves the highest AUROC of 0.922 which shows that this classifier produces confidence scores that best differentiate indicative and ambiguous file names. In comparison, Table 7 displays that all models output higher AUROC for the indicative and out-of-scope confidence scores, further implying that the Random Forest file name classifier outputs confidence scores that can distinguish most indicative file names from all others. The differences in AUROC for different tokenizers are also not substantial for all models.
Figure 3: The confidence score distribution for Random Forest classifier with trie tokenizer on indicative and ambiguous file names from the Common Crawl dataset.
@@ 6.2 Detection Ability for Ambiguous and Out-of-Scope File Names
6.3 Inference Time
Tokenizer SVM Naive Bayes Random Forest XGBoost BERT
trie 0.905 0.917 0.923 0.906 0.876
BERT 0.909 0.915 0.925 0.912 0.886
Table 7: The AUROC of confidence scores between the indicative and out-of-scope file names for all the classifiers with a trie or BERT tokenizer.
Among all the models, Naive Bayes was the fastest, with a average log inference time of $-13.34$, requiring $1.61\times 10^{-6}$ seconds per prediction. The BERT file name classifier’s inference time is substantially slower than all four other models, requiring $6.14\times 10^{-2}$ seconds per prediction. The Random Forest classifier that achieves the highest prediction accuracy on both the Web Search dataset and the Common Crawl dataset requires $1.30\times 10^{-4}$ seconds per prediction.
In comparison, when used on the same computer as the file name classifiers, the BERT model that uses the first 512 tokens of a document content as input requires $0.0678$ seconds per document. Meanwhile, the DiT model on average requires $0.0574$ seconds per image running on a GPU. The robust Random Forest file name classifier is 442x faster than the DiT model even if the DiT model only processes the first page of every document and is 552x faster than the BERT model that processes the content of the documents, offering a dramatic speedup.
@@ 6.3 Inference Time
Figure 4: Overall Accuracy versus Total Inference Time (seconds) per file name at all confidence score threshold between 0 and 1, in increments of 0.001, for a Random Forest classifier with the trie tokenizer predicting on all the file names in the Common Crawl dataset, with a DiT model processing the deferred documents.
6.4 Accuracy and Cost Tradeoffs
Due to the differences in how we obtain confidence scores for each models, the optimal confidence threshold differs empirically between models. Increasing the confidence threshold increases the prediction accuracy while decreasing the prediction rate. In practice, we must select an appropriate confidence threshold to balance the tradeoff between the prediction accuracy and the fraction of documents accurately predicted as indicative — the documents that are deferred as ambiguous or out-of-scope would require further examination using a deeper and more complex model. Thus, the average time and resource expenditure per document classified would increase with the increase in confidence threshold for predicting a document file name as indicative.
Figure 4 shows the accuracy and time inference time per document using the Random Forest file name classifier predicting on all the file names in the Common Crawl dataset, and DiT model processing the deferred documents. With the assumption that all DiT predictions are correct, accuracy is calculated by prediction accuracy $\times$ prediction rate $+$ ($1-$ prediction rate), and total time is calculated by average inference time of Random Forest + average inference time of DiT * ($1-$ prediction rate). The inference time for each model is discussed in Section 6.3. At a confidence threshold of 0.31, the improvement in accuracy begins to taper off, while the total time increases more steeply, indicating that an optimal confidence threshold should be chosen close to that value during deployment. Due to the Common Crawl dataset being randomly sampled from the web without attempting to match the label space of the Web Search dataset, only 23.7% file names in the Common Crawl dataset are indicative and in-scope. Since the majority of the file names are ambiguous or out-of-scope, the file name classifier defers those predictions, which leads to a smaller speedup illustrated in the figure. Thus, better alignment between the label spaces of the training and testing data would lead to more drastic speedups in end-to-end performance.
6.5 Failure Modes
We analyze the failure modes of the Random Forest file name classifier with the trie tokenizer. We observe that each class produces close to the same number of errors when tested on the Web Search dataset. When this file name classifier is tested against the Common Crawl dataset with a confidence threshold of 0.31, we observe that 7.55% of the ambiguous file names are incorrectly classified, while 12.55% of the out-of-scope file names are not deferred.
@@ 6.6 Trade-offs for $k$ Value Selection
In Section 4.1, we mentioned that tokens with TF-IDF larger than $k$ are considered “keywords” and added to the tokenizer. We examine the trade-offs for the selection of $k$ value for the trie tokenizer by comparing the prediction accuracy and prediction rate with various $k$ values. The previous subsections indicate that the Random Forest file name classifier produces the highest prediction accuracy with a fixed minimum prediction rate. Figure 5 displays the accuracy and prediction for the Random Forest file name classifiers with $k$ values from 0 to 0.3. For $k$ values between 0 and 0.059, the accuracy changes sporadically. We observe that the accuracy increases substantially when $k$ values reaches 0.196, then remains constant as $k$ values increases. In this study, we select the $k$ values to be 0.2.
Figure 5: Accuracy of the Random Forest file name classifier with a trie tokenizer with all k values between 0 and 0.3, incremented by 0.001, tested on the Web Search dataset.
@@ 7 Conclusion
We present a new approach to multi-class document classification that dramatically improves classification efficiency while achieving high accuracy. Additionally, we contribute two new document classification datasets that contain file names. Our evaluation shows that a Random Forest classifier with a trie tokenizer, when tasked to predict at least 90% of the data, can achieve 99.78% prediction accuracy on indicative file names from the Web Search dataset and 96.6 % prediction accuracy on the indicative file names from the Common Crawl dataset. The Random Forest classifier with a trie tokenizer can also produce confidence scores that can effectively distinguish indicative and ambiguous file names, resulting in an AUROC of 0.922 for confidence scores between indicative and ambiguous. Meanwhile, the Random Forest classifier is 442x faster than the DiT model, indicating that uses the Random Forest file name classifier can provide a substantial overall speedup in document classification. We hope that our technique and analysis will enable faster and computationally efficient document classifiers without sacrificing on accuracy.
Limitations
In this study, we solely focus on document classification using file names and obtained promising results. We believe metadata — such as file size, page count, number of images — can be used along with file names to improve the classifier accuracy without substantial increase in inference time. However, due to time constraints, we are not able to explore this approach. Future work can investigate the potential benefits of combining metadata with file name to improve classification accuracy and handle ambiguous data or out-of-scope data more effectively.
@@ Appendix A Appendix
We provide both the Web Search and Common Crawl datasets, and the tokenizers in an anonymous Github repository: https://github.com/frank7li/Document_Classification_using_File_Names.
@@
Table 9: Document categories used in the Web Search dataset.